# Periodic Vibration Gaussian: Dynamic Urban Scene Reconstruction and Real-time Rendering

### v2

@@ Abstract
Modeling dynamic, large-scale urban scenes is challenging due to their highly intricate geometric structures and unconstrained dynamics in both space and time. Prior methods often employ high-level architectural priors, separating static and dynamic elements, resulting in suboptimal capture of their synergistic interactions. To address this challenge, we present a unified representation model, called Periodic Vibration Gaussian (PVG). PVG builds upon the efficient 3D Gaussian splatting technique, originally designed for static scene representation, by introducing periodic vibration-based temporal dynamics. This innovation enables PVG to elegantly and uniformly represent the characteristics of various objects and elements in dynamic urban scenes. To enhance temporally coherent and large scene representation learning with sparse training data, we introduce a novel temporal smoothing mechanism and a position-aware adaptive control strategy respectively. Extensive experiments on Waymo Open Dataset [31] and KITTI benchmarks [13] demonstrate that PVG surpasses state-of-the-art alternatives in both reconstruction and novel view synthesis for both dynamic and static scenes. Notably, PVG achieves this without relying on manually labeled object bounding boxes or expensive optical flow estimation. Moreover, PVG exhibits 900-fold acceleration in rendering over the best alternative.
@@
(b) Remove the dynamic
Figure 1: Our proposed Periodic Vibration Gaussian is crafted to effectively and uniformly capture both the static and dynamic elements of a large, dynamic urban scene. (a) It not only reconstructs a dynamic urban scene but also enables real-time rendering, while efficiently isolating dynamic components from the intricacies of the highly unconstrained and complex scene. (b) This capability facilitates flexible manipulation, such as the removal of dynamic scene elements.
@@ 1 Introduction
To overcome the identified limitations, this paper introduces a novel dynamic scene representation method termed Periodic Vibration Gaussian (PVG). This approach provides a unified representation of both static and dynamic elements within a scene through a single formulation. Building upon the efficient 3D Gaussian Splatting technique [15], originally devised for static scene representation, we incorporate periodic vibration-based temporal dynamics. This modification allows for a cohesive representation of static and dynamic scene elements with explicit motion properties such as velocity and staticness. We also propose a position-aware point adaptive control strategy to fit the distant view better. To improve the temporal continuity in representation learning with typically limited training data, we introduce a novel temporal smoothing mechanism.
Our contributions are summarized as follows: (i) Introduction of the very first unified representation model, Periodic Vibration Gaussian (PVG), for large-scale dynamic urban scene reconstruction. In contrast to previous NeRF-based solutions, PVG employs the 3D Gaussian Splatting paradigm, uniquely extending it to elegantly represent dynamic scenes. This is accomplished by seamlessly integrating periodic vibration-based temporal dynamics into the conventional 3D Gaussian Splatting formulation. (ii) Development of a novel temporal smoothing mechanism to enhance the temporal continuity of representation and a position-aware point adaptive control strategy for unbounded urban scenes. (iii) Extensive experiments on two large benchmarks (KITTI and Waymo) demonstrate that PVG outperforms all previous state-of-the-art alternatives in novel view synthesis. Moreover, it provides significant efficiency benefits in both training and rendering processes, achieving a remarkable 900-fold acceleration in rendering compared to the leading competitor, EmerNeRF [45].
@@ 2 Related work
To tackle these challenges, several studies have proposed solutions to enhance training speed. Techniques such as voxel grids [30], hash encoding [22], and tensor factorization [8, 9] have been explored. Others have focused on improving rendering speed by transforming implicit volumes into explicit textured meshes, as demonstrated in works like [10, 27, 49]. Additionally, endeavors such as [3, 4, 36, 5] aim to enhance rendering quality by addressing issues like antialiasing and reflection modeling. More recently, 3D Gaussian Splatting (3DGS) introduced an innovative point-based 3D scene representation, seamlessly integrating the high-quality volume rendering principles of NeRF [21] with the swift rendering speed characteristic of rasterization.
@@
Dynamic scene models Reconstructing dynamic scenes poses distinctive challenges, particularly in effectively handling temporal correlations across various time steps. Expanding on the accomplishments of NeRF [21], several extensions have been proposed to tailor NeRF to dynamic scenarios. In one research direction, certain studies [18, 11, 7, 38, 2] introduce time as an additional input to the radiance field, treating the scene as a 6D plenoptic function. However, this approach couples positional variations induced by temporal dynamics with the radiance field, lacking geometric priors about how time influences the scene. An alternative approach [26, 24, 25, 33, 1, 20, 40] focuses on modeling the movement or deformation of specific static structures, assuming that the dynamics arise from these static elements within the scene. Point-based methods [1, 20, 40] have shown promise in addressing the challenges of reconstructing dynamic scenes due to their flexibility. Building upon the progress in 3D Gaussian splatting, recent works [48, 40] propose the use of a set of deformable 3D Gaussians for optimization across different timestamps. While this kind of methods need to learn the deform function in the dense space which is difficult to migrate to large scenarios. [47] extends the 3D gaussian expression to a 4D form leading to confusion between dynamic and static distinctions.
@@
Urban scene reconstruction NeRF-based techniques have shown their efficacy in autonomous driving scenarios [13, 31]. One research avenue has focused on enhancing the modeling of static street scenes by utilizing scalable representations [32, 34, 19, 28], achieving high-fidelity surface reconstruction [28, 39, 14], and incorporating multi-object composition [43]. However, these methods face difficulties in handling dynamic elements commonly encountered in autonomous driving contexts. Another research direction seeks to address these challenges. Notably, these techniques require additional input, such as leveraging panoptic segmentation to refine the dynamics of reconstruction [17]. Moreover, in [23], scene graphs are employed to decompose dynamic multi-object scenes, while in [46], neural shape priors are learned for completing dynamic object reconstructions. In [41] foreground instances and background environments are decomposed. [44, 50] decompose the scene with different set of gaussian points by bounding boxes. However, they all need manually annotated or predicted bounding boxes and have difficulty to reconstruct the non-rigid objects. In [35], a scalable hash table is proposed for large-scale dynamic scenes, relying on an off-the-shelf 2D optical flow estimator to track dynamic actors. [45] reduces dependence on optical flow by self-supervision, however, it still suffers low image quality and rendering speed.
@@ 3 Method
Utilizing the sequentially acquired and calibrated multi-sensor data, encompassing a set of images $\mathcal{I}$, each captured by cameras equipped with corresponding intrinsic matrices $\mathbf{I}$ and extrinsic matrices $\mathbf{E}$, alongside their respective capture timestamps $t$, collectively represented as $\{\mathcal{I}_{i},t_{i},\mathbf{E}_{i},\mathbf{I}_{i}|i=1,2,\dots N_{c}\}$, and the spatial coordinates of LiDAR point clouds annotated with timestamps $\{(x_{i},y_{i},z_{i},t_{i})|i=1,2,\dots N_{l}\}$, where $N_{c}$ and $N_{l}$ are the number of image frames and Lidar points, our research aims to achieve precise 3D reconstruction and synthesize novel viewpoints at any desired timestamp $t$ and camera pose $[\mathbf{E}_{o},\mathbf{I}_{o}]$. To this end, our framework is meticulously engineered to approximate a rendering function ^$\hat{\mathcal{I}}=\mathcal{F}_{\theta}(\mathbf{E}_{o},\mathbf{I}_{o},t)$.
@@ 3.1 Preliminary
3D Gaussian Splatting (3DGS) [15] utilizes a collection of 3D Gaussians to represent a scene. Through a tile-based rasterization process, 3DGS facilitates real-time alpha blending of numerous Gaussians. The scene is modeled by a set of points $\{P_{i}\}$, where each point $P$ is linked to a mean $\bm{\mu}\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, an opacity $o$, and a color $\mathbf{c}$. These attributes collectively define the point’s influence within the 3D space as:
@@
where $\alpha$ is derived through the product of the opacity ${o}$ and the contribution from the 2D covariance calculated using $\Sigma^{\prime}$ and the corresponding pixel coordinates in the image space. The covariance matrix holds a meaningful interpretation when it is positive semi-definite. In the context of 3DGS, it is decomposed into a scaling matrix, denoted as a diagonal matrix represented by $\bm{s}\in\mathbb{R}^{3}$, and a rotation matrix represented by a unit quaternion $\bm{q}$.
@@
^$\hat{\mathcal{I}}=\mathrm{Render}(\{\mathcal{C}_{i}|i=1,2,\dots,N\};\mathbf{E},\mathbf{I}),$ (4)
@@
Limitation The 3D Gaussian Splatting model represents static points in a scene, lacking the ability to capture dynamic changes over time, essential for modeling dynamic urban scenes. To address this limitation, we propose the Periodic Vibration Gaussian model.
@@ 3.2 Periodic Vibration Gaussian (PVG)
Figure 2: (a) Dynamic and static can be expressed uniformly through different Periodic Vibration Gaussian points with different lifespan. Our model can be trained to differentiate between dynamic and static on its own. (b) Our training pipeline.
@@
$\mathcal{H}(t)$ =$=\{\widetilde{\bm{\mu}}(t),\bm{q},\bm{s},\widetilde{o}(t),\bm{c}\},$ (5)
@@
Conversely, the static components of a scene can be effectively represented by points exhibiting large $\rho$. Introducing a threshold on $\rho$ enables us to discern whether a point represents dynamic elements (see Fig. 1(b)).
@@
^$\hat{\mathcal{I}}=\mathrm{Render}(\{\mathcal{H}_{i}(t)|i=1,\dots,N_{H}\};\mathbf{E},\mathbf{I}),$ (8)
@@ 3.3 Position-aware point adaptive control
$\begin{split}\gamma(\bm{\mu})=\left\{\begin{array}[]{ll}1&\mathrm{if}\quad\|\bm{\mu}\|_{2}<2r\\ \|\bm{\mu}\|_{2}/r-1&\mathrm{if}\quad\|\bm{\mu}\|_{2}\geq 2r,\end{array}\right.\end{split}$ (9)
@@ 3.4 Model training
The intuition comes from the average velocity weighted by opacity decay which is hard to compute the explicit solution, i.e.,
@@
^$\widehat{\mathcal{H}}(t_{2})=\{\widetilde{\bm{\mu}}(t_{1})+\bm{\bar{v}}\cdot\Delta t,\bm{q},\bm{s},\widetilde{o}(t_{1}),\bm{c}\}.$ (12)
@@
Sky refinement To handle the high-frequency details in the sky, we utilize a high-resolution learnable environment cube map $f_{sky}(d)=c_{sky}$ as the background. The final color is articulated as $C_{f}=C+(1-O)f_{sky}(d)$, where $O=\sum^{N_{H}}_{i=1}T_{i}\alpha_{i}$ represents the rendered opacity. During the training phase, we incorporate random perturbations to the ray direction $d$ within its unit pixel length to enhance anti-aliasing.
@@ 4 Experiments
(a) 3DGS [15]
(b) SUDS [35]
(c) EmerNeRF [45]
(d) PVG (Ours)
(e) GT
Figure 4: Novel view synthesis of dynamic scenes on Waymo.
@@
Competitors For dynamic scenes, we evaluate our method alongside S-NeRF [43], StreetSurf [14], Mars [41], 3DGS [15], NSG [23], SUDS [35] and EmerNeRF [45]. In line with previous methods, we conduct evaluations on both image reconstruction and novel view synthesis tasks, selecting every fourth timestamp from each camera as the test set for novel view synthesis. Although our primary focus is on dynamic scenes, to be fair, we also provide a quantitative comparison with methods [43, 14, 15] tailored for static scenes. This is to demonstrate our model’s superiority at uniformly managing both static and dynamic environments.
@@ 4.1 Comparison with state of the art
Table 1: Quantitative results on dynamic scenes on Waymo and KITTI.
Waymo Open Dataset KITTI
Image reconstruction Novel view synthesis Image reconstruction Novel view synthesis
@@
S-NeRF [43] 0.0014 19.67 0.528 0.387 19.22 0.515 0.400 0.0075 19.23 0.664 0.193 18.71 0.606 0.352
StreetSurf [14] 0.097 26.70 0.846 0.3717 23.78 0.822 0.401 0.37 24.14 0.819 0.257 22.48 0.763 0.304
3DGS [15] 63 27.99 0.866 0.293 25.08 0.822 0.319 125 21.02 0.811 0.202 19.54 0.776 0.224
NSG [23] 0.032 24.08 0.656 0.441 21.01 0.571 0.487 0.19 26.66 0.806 0.186 21.53 0.673 0.254
Mars [41] 0.030 21.81 0.681 0.430 20.69 0.636 0.453 0.31 27.96 0.900 0.185 24.23 0.845 0.160
SUDS [35] 0.008 28.83 0.805 0.289 21.83 0.656 0.405 0.04 28.31 0.876 0.185 22.77 0.797 0.171
EmerNeRF [45] 0.053 28.11 0.786 0.373 25.92 0.763 0.384 0.28 26.95 0.828 0.218 25.24 0.801 0.237
PVG (Ours) 50 32.46 0.910 0.229 28.11 0.849 0.279 59 32.83 0.937 0.070 27.43 0.896 0.114
PVG (5cam) 41 32.05 0.895 0.252 27.50 0.828 0.312 - - - - - - -
@@
In our experiments, we utilize the three frontal cameras to assess performance on four challenging dynamical scenes (each contains around 50 frames), chosen due to their substantial movement. Table 1 displays the average error metrics across these selected dynamic scenes for both image reconstruction and novel view synthesis tasks. Our model markedly outperforms baselines [43, 14, 15, 23, 35, 41] across all metrics in both tasks. Specifically, for the image reconstruction task, we note a $12.6\%$ increase in PSNR, $13.0\%$ in SSIM, and a $20.8\%$ decrease in LPIPS compared to the leading SUDS [35] baseline. For novel view synthesis, our technique synthesizes high-quality views of unseen timestamps and also significantly surpasses the best EmerNeRF [45] performance by $8.4\%$ in PSNR, $11.3\%$ in SSIM, and $27.3\%$ in LPIPS. These enhancements are visually corroborated in Fig. 4, where our method faithfully captures both static and dynamic elements with remarkable detail. Noteworthy is our method’s efficiency, completing the entire training process around an hour, and stands out as the method with the fastest training speed compared to the baselines. We also report training results of our method on fully 5 cameras, for we use every forth frame of 5 cameras as our testing set, our performance dropped off slightly. Our rendering speed, measured in Frames Per Second (FPS), significantly outperforms competing methods and stands slightly below that of 3DGS [15].
For fair comparison with EmerNeRF [45], we adopt its training setup, i.e., 4$\times$ downsample rate for image resolution and using the whole sequences for training. We randomly selected four Waymo scenes for test in Table 3. It is shown that PVG is clearly superior over EmerNeRF. More visualization is given in Fig. 8.
Table 2: Dynamic scenes’ PSNR, same settings as EmerNeRF [45].
Table 3: Static scenes’ PSNR, same settings as StreetSurf [14].
Sequence Image Reconstruction Novel View Synthesis
EmerNeRF PVG (Ours) EmerNeRF PVG (Ours)
seg102319… 28.93 34.15 28.64 29.52
seg117188… 28.42 32.34 26.29 27.16
seg103913… 29.10 32.01 27.73 28.47
seg109636… 28.39 31.46 27.35 28.28
Average 28.69 32.49 27.50 28.35
Sequence Image Reconstruction
StreetSurf PVG (Ours)
seg153495… 27.26 31.22
seg405841… 28.08 34.91
seg342571… 29.42 32.37
seg134763… 28.20 31.11
Average 28.24 32.40
Table 3: Static scenes’ PSNR, same settings as StreetSurf [14].
@@
Figure 6: Novel view synthesis on static scene on Waymo.
@@
Figure 7: PVG vs. EmerNeRF [45].
Figure 8: Rendered rgb, depth and semantic label.
@@
We also show that by making use of the 2D semantic labels provided in a video sequence, our method can derive semantic categories in novel viewpoints or timestamps and render as shown in Fig. 8. This is achieved by assigning a 19-dimension vector for each Gaussian, representing the probability of which category the Gaussian belongs to and the rendered probability map is supervised by 2D semantic labels with cross-entropy.

### v3

@@ Abstract
Modeling dynamic, large-scale urban scenes is challenging due to their highly intricate geometric structures and unconstrained dynamics in both space and time. Prior methods often employ high-level architectural priors, separating static and dynamic elements, resulting in suboptimal capture of their synergistic interactions. To address this challenge, we present a unified representation model, called Periodic Vibration Gaussian (PVG). PVG builds upon the efficient 3D Gaussian splatting technique, originally designed for static scene representation, by introducing periodic vibration-based temporal dynamics. This innovation enables PVG to elegantly and uniformly represent the characteristics of various objects and elements in dynamic urban scenes. To enhance temporally coherent and large scene representation learning with sparse training data, we introduce a novel temporal smoothing mechanism and a position-aware adaptive control strategy respectively. Extensive experiments on Waymo Open Dataset [1] and KITTI benchmarks [2] demonstrate that PVG surpasses state-of-the-art alternatives in both reconstruction and novel view synthesis for both dynamic and static scenes. Notably, PVG achieves this without relying on manually labeled object bounding boxes or expensive optical flow estimation. Moreover, PVG exhibits 900-fold acceleration in rendering over the best alternative. The code is available at https://github.com/fudan-zvg/PVG.
keywords:
Dynamic Urban Scene, 3D Reconstruction, Gaussian Splatting
@@
(b) Remove the dynamic scene elements
Figure 1: Our proposed Periodic Vibration Gaussian is crafted to effectively and uniformly capture both static and dynamic elements of a large, dynamic urban scene. (a) It not only reconstructs a dynamic urban scene but also enables real-time rendering, while efficiently isolating dynamic components from the intricacies of the highly unconstrained and complex scene. (b) This capability facilitates flexible manipulation, such as the removal of dynamic scene elements.
@@ 1 Introduction
To overcome the identified limitations, this paper introduces a novel dynamic scene representation method termed Periodic Vibration Gaussian (PVG). This approach provides a unified representation of both static and dynamic elements within a scene through a single formulation. Building upon the efficient 3D Gaussian Splatting (3DGS)) [11], originally devised for static scene representation, we incorporate periodic vibration-based temporal dynamics. This modification allows for a cohesive representation of static and dynamic scene elements with explicit motion properties such as velocity and staticness. We also propose a position-aware point adaptive control strategy to fit the distant view better. To improve the temporal continuity in representation learning with typically limited training data, we introduce a novel temporal smoothing mechanism.
Our contributions are summarized as follows: (i) Introduction of the very first unified representation model, PVG, for large-scale dynamic urban scene reconstruction. In contrast to previous NeRF-based solutions, PVG employs the 3D Gaussian Splatting paradigm, uniquely extending it to elegantly represent dynamic scenes. This is accomplished by seamlessly integrating periodic vibration-based temporal dynamics into the conventional 3DGS formulation. (ii) Development of a novel temporal smoothing mechanism to enhance the temporal continuity of representation and a position-aware point adaptive control strategy for unbounded urban scenes. (iii) Extensive experiments on two large benchmarks (KITTI and Waymo) demonstrate that PVG outperforms all previous state-of-the-art alternatives in novel view synthesis. Moreover, it provides significant efficiency benefits in both training and rendering processes, achieving a remarkable 900-fold acceleration in rendering compared to the leading competitor, EmerNeRF [10]. We also show that PVG is superior in both visual quality and rendering efficiency over concurrent 3DGS based models.
@@ 2 Related work
To tackle these challenges, several studies have proposed solutions to enhance training speed. Techniques such as voxel grids [13], hash encoding [14], and tensor factorization [15, 16] have been explored. Others have focused on improving rendering speed by transforming implicit volumes into explicit textured meshes, as demonstrated in works like [17, 18, 19]. Additionally, endeavors such as [20, 21, 22, 23] aim to enhance rendering quality by addressing issues like antialiasing and reflection modeling. Recently, 3D Gaussian Splatting (3DGS) [11] introduces an innovative point-based 3D scene representation, seamlessly integrating the high-quality volume rendering principles of NeRF with the swift rendering speed characteristic of rasterization.
@@
Dynamic scene models Reconstructing dynamic scenes poses distinctive challenges, particularly in effectively handling temporal correlations across various time steps. Expanding on the accomplishments of NeRF [12], several extensions have been proposed to tailor NeRF to dynamic scenarios. In one research direction, certain studies [24, 25, 26, 27, 28] introduce time as an additional input to the radiance field, treating the scene as a 6D plenoptic function. However, this approach couples positional variations induced by temporal dynamics with the radiance field, lacking geometric priors about how time influences the scene. An alternative approach [29, 30, 31, 32, 33, 34, 35] focuses on modeling the movement or deformation of specific static structures, assuming that the dynamics arise from these static elements within the scene. Point-based methods [33, 34, 35] have shown promise in addressing the challenges of reconstructing dynamic scenes due to their flexibility. Building upon the progress in 3DGS, recent works [36, 35] propose the use of a set of deformable 3D Gaussians for optimization across different timestamps. While this kind of methods need to learn the deform function in the dense space which is difficult to migrate to large scenarios. [37] extends 3DGS to a 4D formulation, enabling modeling of the full time-space manifold. While the formulation induces a dynamic opacity model through spatio-temporal 4D Gaussians, it does not explicitly model the separation between static and dynamic elements. In contrast, we introduce a staticity coefficient to represent the per-point degree of motion, enabling effective disentanglement of static and dynamic components. To address the challenge of temporally sparse observations, we propose a temporal smoothing training strategy that enhances temporal consistency and reconstruction robustness. Furthermore, our framework incorporates several domain-specific designs for autonomous driving scenarios, including LiDAR-projected depth supervision, cube map-based sky modeling, and a positional-aware control mechanism for handling large-scale distant structures. These innovations collectively contribute to higher-quality reconstruction and synthesis in complex urban environments.
@@
Urban scene reconstruction NeRF-based techniques have shown their efficacy in autonomous driving scenarios [2, 1]. One research avenue has focused on enhancing the modeling of static street scenes by utilizing scalable representations [38, 39, 40, 4], achieving high-fidelity surface reconstruction [4, 41, 6], and incorporating multi-object composition [5]. However, these methods face difficulties in handling dynamic elements commonly encountered in autonomous driving contexts. Another research direction seeks to address these challenges. Notably, these techniques require additional input, such as leveraging panoptic segmentation to refine the dynamics of reconstruction [8]. Moreover, in [7], scene graphs are employed to decompose dynamic multi-object scenes, while in [42], neural shape priors are learned for completing dynamic object reconstructions. In [43] foreground instances and background environments are decomposed. 3DGS-based methods [44, 45, 46, 47] have been concurrently proposed for dynamic urban scene reconstruction. However, these approaches generally depend on object-level supervision via 3D bounding boxes. While such supervision can improve the reconstruction quality of dynamic objects, it increases model complexity and reduces flexibility. Moreover, reliance on automatically generated bounding boxes introduces noisy supervision and often fails in challenging cases such as distant or occluded objects. In addition, explicitly separating dynamic objects from the background may compromise the overall consistency of the reconstructed scene. In [9], a scalable hash table is proposed for large-scale dynamic scenes, relying on an off-the-shelf 2D optical flow estimator to track dynamic actors. [10] reduces dependence on optical flow by self-supervision, however, it still suffers low image quality and rendering speed.
@@ 3 Method
Utilizing the sequentially acquired and calibrated multi-sensor data, encompassing a set of images $\mathcal{I}$, each captured by cameras equipped with corresponding intrinsic matrices $\mathbf{I}$ and extrinsic matrices $\mathbf{E}$, alongside their respective capture timestamps $t$, collectively represented as $\{\mathcal{I}_{i},t_{i},\mathbf{E}_{i},\mathbf{I}_{i}|i=1,2,\dots N_{c}\}$, and the spatial coordinates of LiDAR point clouds annotated with timestamps $\{(x_{i},y_{i},z_{i},t_{i})|i=1,2,\dots N_{l}\}$, where $N_{c}$ and $N_{l}$ are the number of image frames and Lidar points, our research aims to achieve precise 3D reconstruction and synthesize novel viewpoints at any desired timestamp $t$ and camera pose $\mathbf{E}_{o},\mathbf{I}_{o}]$. To this end, our framework is meticulously engineered to approximate a rendering function $\hat{\mathcal{I}}=\mathcal{F}_{\theta}(\mathbf{E}_{o},\mathbf{I}_{o},t)$.
@@ 3.1 Preliminary
3DGS [11] utilizes a collection of 3D Gaussians to represent a scene. Through a tile-based rasterization process, 3DGS facilitates real-time alpha blending of numerous Gaussians. The scene is modeled by a set of points $\{P_{i}\}$, where each point $P$ is linked to a mean $\bm{\mu}\in\mathbb{R}^{3}$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, an opacity $o$, and a color $\mathbf{c}$. These attributes collectively define the point’s influence within the 3D space as:
@@
where $\alpha$ is derived through the product of the opacity o{o} and the contribution from the 2D covariance calculated using $\Sigma^{\prime}$ and the corresponding pixel coordinates in the image space. The covariance matrix holds a meaningful interpretation when it is positive semi-definite. In the context of 3DGS, it is decomposed into a scaling matrix, denoted as a diagonal matrix represented by $\bm{s}\in\mathbb{R}^{3}$, and a rotation matrix represented by a unit quaternion $\bm{q}$.
@@
$\hat{\mathcal{I}}=\mathrm{Render}(\{\mathcal{C}_{i}|i=1,2,\dots,N\};\mathbf{E},\mathbf{I}),$ (4)
@@
Limitation The 3DGS model represents static points in a scene, lacking the ability to capture dynamic changes over time, essential for modeling dynamic urban scenes. To address this limitation, we propose the Periodic Vibration Gaussian (PVG) model.
@@ 3.2 Periodic Vibration Gaussian (PVG)
Figure 2: PVG learns an adaptive opacity decay rate to distinguish between dynamic and static scene elements. The model represents dynamic objects with short-lifespan points that quickly fade, while static regions are modeled by points with a longer lifespan, allowing them to exhibit globally consistent behavior over time. This learning process is guided by supervision from RGB and LiDAR depth signals at each time step.
@@
$\mathcal{H}(t)$ $=\{\widetilde{\bm{\mu}}(t),\bm{q},\bm{s},\widetilde{o}(t),\bm{c}\},$ (5)
@@
Conversely, the static components of a scene can be effectively represented by points exhibiting large $\rho$. Introducing a threshold on $\rho$ enables us to discern whether a point represents dynamic elements (see Fig. 1(b)). Conceptually, for the periodic-function-based design, this does not imply that the underlying motion in the scene is inherently periodic; Rather, our periodic formulation can be regarded as a primitive for fitting complex motions, while still preserving the global static property.
@@
$\hat{\mathcal{I}}=\mathrm{Render}(\{\mathcal{H}_{i}(t)|i=1,\dots,N_{H}\};\mathbf{E},\mathbf{I}),$ (8)
@@ 3.3 Position-aware point adaptive control
[…]
@@ 3.4 Model training
The intuition comes from the average velocity weighted by opacity decay
@@
$\widehat{\mathcal{H}}(t_{2})=\{\widetilde{\bm{\mu}}(t_{1})+\bm{\bar{v}}\cdot\Delta t,\bm{q},\bm{s},\widetilde{o}(t_{1}),\bm{c}\}.$ (12)
@@
Sky refinement Representing the static sky using Gaussian points is theoretically possible, but in practice, this would require placing them at extremely large distances with extremely large scales, which causes optimization challenge. To address this, we adopt a cube map representation, where the sky color depends solely on the viewing direction. This approach is physically reasonable, lightweight, and does not affect rendering speed. Specifically, we utilize a high-resolution learnable environment cube map $f_{sky}(d)=c_{sky}$ as the background. The final color is articulated as $C_{f}=C+(1-O)f_{sky}(d)$, where $O=\sum^{N_{H}}_{i=1}T_{i}\alpha_{i}$ represents the rendered opacity. During the training phase, we incorporate random perturbations to the ray direction $d$ within its unit pixel length to enhance anti-aliasing.
@@ 4 Experiments
Competitors For dynamic scenes, we evaluate our method alongside S-NeRF [5], StreetSurf [6], Mars [43], 3DGS [11], NSG [7], SUDS [9] and EmerNeRF [10]. In line with previous methods, we conduct evaluations on both image reconstruction and novel view synthesis tasks, selecting every fourth timestamp from each camera as the test set for novel view synthesis. Although our primary focus is on dynamic scenes, to be fair, we also provide a quantitative comparison with methods [5, 6, 11] tailored for static scenes. This is to demonstrate our model’s superiority at uniformly managing both static and dynamic environments. Furthermore, we also compare PVG with concurrent approaches [44, 45, 49, 47, 37] based on bounding boxes or 3D Gaussian Splatting. The comparison includes both the reconstruction quality of the static background and the dynamic foreground.
@@ 4.1 Comparison with state of the art
Table 1: Quantitative results on dynamic scenes on Waymo and KITTI. The bold text denotes the best result.
Waymo Open Dataset
Image reconstruction Novel view synthesis
@@
S-NeRF [5] 0.0014 19.67 0.528 0.387 19.22 0.515 0.400
StreetSurf [6] 0.097 26.70 0.846 0.3717 23.78 0.822 0.401
3DGS [11] 63 27.99 0.866 0.293 25.08 0.822 0.319
NSG [7] 0.032 24.08 0.656 0.441 21.01 0.571 0.487
Mars [43] 0.030 21.81 0.681 0.430 20.69 0.636 0.453
SUDS [9] 0.008 28.83 0.805 0.289 21.83 0.656 0.405
EmerNeRF [10] 0.053 28.11 0.786 0.373 25.92 0.763 0.384
PVG (Ours) 50 32.46 0.910 0.229 28.11 0.849 0.279
PVG (5cam) 41 32.05 0.895 0.252 27.50 0.828 0.312
KITTI
Image reconstruction Novel view synthesis
[…]
S-NeRF [5] 0.0075 19.23 0.664 0.193 18.71 0.606 0.352
StreetSurf [6] 0.37 24.14 0.819 0.257 22.48 0.763 0.304
3DGS [11] 125 21.02 0.811 0.202 19.54 0.776 0.224
NSG [7] 0.19 26.66 0.806 0.186 21.53 0.673 0.254
Mars [43] 0.31 27.96 0.900 0.185 24.23 0.845 0.160
SUDS [9] 0.04 28.31 0.876 0.185 22.77 0.797 0.171
EmerNeRF [10] 0.28 26.95 0.828 0.218 25.24 0.801 0.237
PVG (Ours) 59 32.83 0.937 0.070 27.43 0.896 0.114
@@
In our experiments, we utilize the three frontal cameras to assess performance on four challenging dynamical scenes (each contains around 50 frames), chosen due to their substantial movement. Table 1 displays the average error metrics across these selected dynamic scenes for both image reconstruction and novel view synthesis tasks. Our model markedly outperforms baselines [5, 6, 11, 7, 9, 43] across all metrics in both tasks. Specifically, for the image reconstruction task, we note a $12.6\%$ increase in PSNR, $13.0\%$ in SSIM, and a $20.8\%$ decrease in LPIPS compared to the leading SUDS [9] baseline. For novel view synthesis, our technique synthesizes high-quality views of unseen timestamps and also significantly surpasses the best EmerNeRF [10] performance by $8.4\%$ in PSNR, $11.3\%$ in SSIM, and $27.3\%$ in LPIPS. We show more visualization in Fig. 19 and 20. Noteworthy is our method’s efficiency, completing the entire training process around an hour, and stands out as the method with the fastest training speed compared to the baselines. We also report training results of our method on fully 5 cameras, for we use every forth frame of 5 cameras as our testing set, our performance dropped off slightly. Our rendering speed, measured in Frames Per Second (FPS), significantly outperforms competing methods and stands slightly below that of 3DGS [11].
[…]
@@
Figure 5: Novel view synthesis of a static scene on Waymo.
Table 2: Dynamic scenes’ PSNR, same settings as EmerNeRF [10].
Sequence Image Reconstruction Novel View Synthesis
EmerNeRF PVG (Ours) EmerNeRF PVG (Ours)
seg1023… 28.93 34.15 28.64 29.52
seg1171… 28.42 32.34 26.29 27.16
seg1039… 29.10 32.01 27.73 28.47
seg1096… 28.39 31.46 27.35 28.28
Average 28.69 32.49 27.50 28.35
Table 3: Static scenes’ PSNR, same settings as StreetSurf [6].
Sequence Image Reconstruction
StreetSurf PVG (Ours)
seg1534… 27.26 31.22
seg4058… 28.08 34.91
seg3425… 29.42 32.37
seg1347… 28.20 31.11
Average 28.24 32.40
@@
Figure 7: Rendered RGB image, depth and semantic label.
@@
We show that by making use of the 2D semantic labels provided in a video sequence, our method can derive semantic categories in novel viewpoints or timestamps and render as shown in Fig. 7. This is achieved by assigning a 19-dimension vector for each Gaussian, representing the probability of which category the Gaussian belongs to and the rendered probability map is supervised by 2D semantic labels with cross-entropy.